# Active Learning for Control-Oriented Identification of Nonlinear Systems

### v1

@@ A.2 Improved Rates via the Delta Method
The superscript $\pi$ denotes the policy that is rolled out to generate $U_{t}^{k}$. Given sufficiently many trajectories, this estimate converges to the expected value $\Sigma^{\pi}$ (defined in (7)). As long as this covariance matrix is positive definite, $\Sigma^{\pi}\succ 0$, we can show that the delta method is successful in improving the estimation rate.
In order to do so, we require sufficiently many trajectories. In addition to the burn-in for the excess risk bound, we require that $N$ is large enough that Corollary A.1 guarantees $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ with high probability. We refer to the burn-in required for this to hold as $K\geq\tau_{\mathsf{small\,error}}(\delta)$, where
$\tau_{\mathsf{small\,error}}(\delta)=\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right).$
Furthermore, we require that the higher order terms ($\mathsf{h.o.t.}$) from the taylor expansion involved in the Delta method are dominated by the lower order terms. For downstream use of the delta method, we allow this to depend on the norm of an arbitrary positive definite matrix $H$. We refer to the required burn-in as $K\geq\tau_{\mathsf{h.o.t.}}(\delta,H)$, where
$\tau_{\mathsf{h.o.t}}(\delta)$ $=\Bigg{(}\frac{8T^{1-\alpha}L_{f}^{2}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
$+\frac{8d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\sqrt{32T\log\frac{4}{\delta}}\Bigg{)}^{\frac{1}{\alpha}}.$
@@
Let Assumption 2.1 hold and fix $\delta\in(0,\frac{1}{2}]$. Let $H$ be a positive semi-definite matrix. Fix some $\delta\in(0,\frac{1}{4}]$. Consider setting $\hat{\phi}$ as the least squares solution from $K$ episodes of data collected with policy $\pi$ for
$K$ $\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/2),\tau_{\mathsf{small\,error}}(\delta),\tau_{\mathsf{h.o.t}}(\delta,H)\right\}\mathclose{}$
@@
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}$.
As long as $\Sigma^{\pi}\succ 0$ and $K$ is sufficiently large that $\Sigma_{K}^{\pi}$ concentrates to $\Sigma^{\pi}$, we may use the above result to show that $\hat{\phi}_{K}-\phi^{\star}$ behaves like a self-normalized martingale. We may then invoke Theorem 7.2 of Wagenmaker et al. (2021) to bound this self-normalized martingale, leading to Theorem 3.1. Before proving the above lemma, we introduce several preliminary results.
@@
Let $G=-A^{-1}U(\Sigma^{-1}+V^{\top}A^{-1}U)^{-1}V$. By submultiiplicativity, $\left\|G\right\|_{\mathsf{op}}\leq\frac{1}{\lambda_{\min}(A)}\left(\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)\right)^{-1}.$ We have that
@@
By the fact that $\hat{\phi}_{K}$ solves (3),
@@
A second order taylor expansion of $\psi(\hat{\phi}_{K})$ about $\phi^{\star}$ provides
$\psi(\hat{\phi}_{K})$ $=\psi(\phi^{\star})+D_{\phi}\psi(\phi)|_{\phi=\phi_{\star}}(\hat{\phi}_{K}-\phi^{\star})+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}],$
for some $\tilde{\phi}$ which is a convex combination of $\phi^{\star}$ and $\hat{\phi}_{K}$. Here $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot,\cdot]$ is a tensor mapping $\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}^{d_{\phi}}$. We write $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot]$ as the tensor mapping $\mathbb{R}^{d_{\phi}}\to\mathbb{R}^{d_{\phi}\times d_{\phi}}$ such that
$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}]=D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}](\hat{\phi}_{K}-\phi^{\star}).$
As long as $D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ is non-singular, then the above two equations imply
$\hat{\phi}_{K}$ $-\phi^{\star}=\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star}).$
@@
$D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}$ $=K\hat{\Sigma}_{K}^{\pi}+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{2}f_{t}^{k})[W_{t}^{k}]}_{E_{1}}$
$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ $=\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}]Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})}_{E_{2}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})^{\top}(D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}])^{\top}}_{E_{3}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{3}f(X_{t}^{k},U_{t}^{k};\tilde{\phi}))[\hat{\phi}_{K}-\phi^{\star},W_{t}^{k}]}_{E_{4}}$
We will bound the operator norms of $E_{1},E_{2},E_{3}$ and $E_{4}$ in order to apply Lemma A.6. We may bound $E_{2}$ and $E_{3}$ directly using Assumption 2.1:
@@
By the fact that $K\geq\tau_{\mathbf{Err}}(\delta/2)$, the conditions on Corollary A.1 are satisifed such that with probability at least $21-\delta/$ 2,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}.$
Using the burn-in $K\geq\tau_{\mathsf{small\,error}}(\delta)$, we have that this is at most $1$. Therefore, with probability at least $1-\delta$,
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq 2d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}+2(TK)^{1-\alpha}L_{f}^{2}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
As $K\geq\tau_{\mathsf{h.o.t}}(\delta,H)$, we have that
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left((H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}K\lambda_{\min}\left(\hat{\Sigma}_{K}^{\pi}\right).$
We may therefore invoke Lemma A.6 with $s=\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}$ to show that
$\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star})=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k},$
for some $G$ satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}.$
@@
To proceed with our proof of the main result, we will leverage Lemma A.5. In particular, we require a bound on the self-normalized martingale that arises. We state one such bound due to Wagenmaker et al. (2021) for this purpose. The bound is stated in order to characterize the norm of a self-normalized martingale with respect to an arbitrary positive definite matrix $H$.
Lemma A.8 (Theorem 7.2 of Wagenmaker et al. (2021)).
Suppose $\Sigma^{\pi}\succ 0$ and $H\succeq 0$. Define the event $\mathcal{E}\triangleq\mathopen{}\left\{\left\|\Sigma_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}\right\}\mathclose{})$ for some $\beta\in(0,\frac{1}{4})$. As long as $\mathcal{E}$ holds, the following holds with probability at least $1-\delta$,
$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 5(1+\alpha)\sigma_{w}^{2}\log\frac{6d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma^{\pi}\right)^{-1}\right)}{K},$
where $\alpha=26\beta^{2}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right))$.
@@
The proof follows from the proof of Theorem 7.2 of Wagenmaker et al. (2021) by making the following substitutions for $\bm{Z}$ and $\bm{w}$ in the proof:
@@
To guarantee concentration of the empirical covariance (30) to the true covariance (7), we state the following covariance concentration result.
@@
Let $\beta\in(0,\frac{1}{4})$. Suppose
$K\geq\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta)}{\beta^{2}\lambda_{\min}(\Sigma^{\pi})^{2}}.$
@@
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}).$

### v2

@@ A.2 Improved Rates via the Delta Method
The superscript $\pi$ denotes the policy that is rolled out to generate $U_{t}^{k}$. Given sufficiently many trajectories, this estimate converges to the expected value
$\Sigma^{\pi}\triangleq\mathbb{E}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right].$ (26)
As long as this covariance matrix is positive definite, $\Sigma^{\pi}\succ 0$, we can show that the delta method is successful in improving the estimation rate.
In order to do so, we require sufficiently many trajectories. In addition to the burn-in for the excess risk bound, we require that $K$ is large enough that Corollary A.1 guarantees $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ with high probability. We refer to the burn-in required for this to hold as $K\geq\tau_{\mathsf{small\,error}}(\delta)$, where
$\tau_{\mathsf{small\,error}}(\delta)=\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right).$
Furthermore, we require that the higher order terms ($\mathsf{h.o.t.}$) from the taylor expansion involved in the Delta method are dominated by the lower order terms. For downstream use of the delta method, we allow this to depend on the norm of an arbitrary positive definite matrix $H$ and an arbitrary positive number $\beta$. We refer to the required burn-in as $K\geq\tau_{\mathsf{h.o.t.}}(\delta,\beta)$, where
$\tau_{\mathsf{h.o.t}}(\delta,H)$ $=\Bigg{(}\frac{8T^{1-\alpha}L_{f}^{2}}{\beta}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}+\frac{8d_{\mathsf{X}}\sigma_{w}L_{f}}{\beta}\sqrt{32T\log\frac{4}{\delta}}\Bigg{)}^{\frac{1}{\alpha}}.$
@@
Let Assumption 2.1 hold and fix $\delta\in(0,\frac{1}{2}]$. Let $H$ be a positive definite matrix and $\beta$ a positive number. Fix some $\delta\in(0,\frac{1}{4}]$. Consider setting $\hat{\phi}$ as the least squares solution from $K$ episodes of data collected with policy $\pi$ for
$K$ $\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/2),\tau_{\mathsf{small\,error}}(\delta),\tau_{\mathsf{h.o.t}}(\delta,\beta)\right\}\mathclose{}$
@@
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$.
As long as $\Sigma^{\pi}\succ 0$ and $K$ is sufficiently large that $\Sigma_{K}^{\pi}$ concentrates to $\Sigma^{\pi}$, we may use the above result to show that $\hat{\phi}_{K}-\phi^{\star}$ behaves like a self-normalized martingale. We may then bound this self-normalized martingale, leading to Theorem 3.1. Before proving the above lemma, we introduce several preliminary results.
@@
Let $G=-A^{-1}U(\Sigma^{-1}+V^{\top}A^{-1}U)^{-1}V$. By submult iplicativity, $\left\|G\right\|_{\mathsf{op}}\leq\frac{1}{\lambda_{\min}(A)}\left(\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)\right)^{-1}.$ We have that
@@
By the fact that $K\geq\tau_{\mathbf{Err}}(\delta/2)$, the conditions on Corollary A.1 are satisfied such that with probability at least $21-\delta/$ 2,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}.$ (27)
Using the assumption $K\geq\tau_{\mathsf{small\,error}}(\delta)$, we have that this is at most $1$. Under this event, and by the assumption that $\left\|\phi^{\star}\right\|\leq B-1$, we have that $\hat{\phi}_{K}$ is is in the interior of $\mathopen{}\left\{\phi:\left\|\phi\right\|\leq B\right\}\mathclose{}$. Thus by the fact that $\hat{\phi}_{K}$ solves (3), it holds that
@@
A first order Taylor expansion of $\psi(\hat{\phi}_{K})$ about $\phi^{\star}$ provides
$\psi(\hat{\phi}_{K})$ $=\psi(\phi^{\star})+D_{\phi}\psi(\phi)|_{\phi=\phi_{\star}}(\hat{\phi}_{K}-\phi^{\star})+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}],$
where $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})$ is a trilinear operator such that the vector $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}]$ is defined by the product of the matrix $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot\hat{\phi}_{K}-\phi_{\star}]$ with $\hat{\phi}_{K}-\phi_{\star}$: $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}]=\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot\hat{\phi}_{K}-\phi_{\star}](\hat{\phi}_{K}-\phi_{\star})$, and the satisfies the operator norm bound below:
$\left\|\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right\|_{\mathsf{op}}$ $\leq\sup_{\ell\in[0,1]}\left\|D_{\phi}^{2}\psi(\phi)|_{\phi=\ell\phi^{\star}+(1-\ell)\hat{\phi}_{K}}[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]\right\|_{\mathsf{op}}.$
As long as $D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]$ is non-singular, then the above two equations imply
$\hat{\phi}_{K}$ $-\phi^{\star}=\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right)^{-1}\psi(\phi^{\star}).$
@@
$D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}$ $=K\hat{\Sigma}_{K}^{\pi}+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{2}f_{t}^{k})[\cdot,\cdot,W_{t}^{k}]}_{E_{1}},$
$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]$ $=\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]^{\top}Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})}_{E_{2}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})^{\top}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]}_{E_{3}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{3}f(X_{t}^{k},U_{t}^{k};\tilde{\phi}))[\cdot,\cdot,W_{t}^{k},\hat{\phi}_{K}-\phi^{\star}]}_{E_{4}}$
We will bound the norms of $E_{1},E_{2},E_{3}$ and $E_{4}$ for $\tilde{\phi}$ an arbitrary convex combination of $\hat{\phi}_{K}$ and $\phi_{\star}$ in order to apply Lemma A.6. We may bound $E_{2}$ and $E_{3}$ directly using Assumption 2.1:
@@
Using the fact that $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ along with the bound in (27) it holds that with probability at least $1-\delta$,
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq 2d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}+2(TK)^{1-\alpha}L_{f}^{2}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4L_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
As $K\geq\tau_{\mathsf{h.o.t}}(\delta,\beta)$ and $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$, it holds that
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq\frac{\beta}{2}K\lambda_{\min}\left(\hat{\Sigma}_{K}^{\pi}\right).$
We may therefore invoke Lemma A.6 with $s=\frac{\beta}{2}$ to show that
$\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right)^{-1}\psi(\phi^{\star})=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k},$
for some $G$ satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$.
@@
To proceed with our proof of the main result, we will leverage Lemma A.5. In particular, we require a bound on the weighted norm of a self-normalized martingale.
Lemma A.8.
Suppose $\Sigma^{\pi}\succ 0$ and $H\succ 0$. Define the event $\mathcal{E}\triangleq\mathopen{}\left\{\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\right\}\mathclose{}$ for some positive $\beta$ satisfying
$\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$
As long as $\mathcal{E}$ holds, the following holds with probability at least $1-\delta$,
$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 2\sigma_{w}^{2}\left(1+\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}\left(1+\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{1}{\delta}.$
@@
Define
@@
such that
$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}=\left\|(\bm{Z}\bm{Z}^{\top})^{-1}\bm{Z}\bm{W}\right\|_{H}^{2}\mbox{ and }\hat{\Sigma}_{K}^{\pi}=\frac{1}{K}\bm{Z}\bm{Z}^{\top}.$
We may write
$\left\|\left(\bm{Z}\bm{Z}^{\top}\right)^{-1}\bm{Z}\bm{W}\right\|_{H}^{2}$ $=\left\|H^{1/2}\left(\bm{Z}\bm{Z}^{\top}\right)^{-1}\bm{Z}\bm{W}\right\|^{2}$
$=\left\|\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
$=\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
$=2\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}+\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
We have by assumption that $\left\|\frac{1}{K}\bm{Z}\bm{Z}^{\top}-\Sigma^{\pi}\right\|\leq\beta$, and therefore that
$H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\succeq\left(1-\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)H^{-1/2}K\Sigma^{\pi}H^{-1/2}.$
It follows from the above inequality and the assumption that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$ that
$2\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}+\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\left\|\left(\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}+\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
$=2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left\|\left(\frac{\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}}{\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)}+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}.$
The above quantity is a self-normalized martingale. In particular, we can bound it by applying the bound in Theorem 4.1 of Ziemann et al. (2023). To do so, define $\tilde{\Sigma}\triangleq\frac{\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}}{\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)}$. Then it follows that with probability at least $1-\delta$ that
$2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left\|\left(\tilde{\Sigma}+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\log\operatorname{\textbf{det}}\left(I+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right),$
where the second inequality follows by noting that for any positive definite matrix matrix $M\in\mathbb{R}^{d\times d}$, $\log\operatorname{\textbf{det}}(I+M)=\sum_{i=1}^{d}\log(1+\lambda_{i}(M))\leq\sum_{i=1}^{d}\lambda_{i}(M)$. By the concentration event on $\bm{Z}\bm{Z}^{\top}$ and the definition of $\tilde{\Sigma}$, this quantity may be bounded as
$2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\Sigma^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\left(1+\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
$=2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\left(\sigma_{w}^{2}\left(1+\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}\left(H\left(K\Sigma^{\pi}\right)^{-1}\right)+2\sigma_{w}^{2}\left\|H\left(K\Sigma^{\pi}\right)^{-1}\right\|\log\frac{1}{\delta}\right)$
$\leq 2\left(1+\beta\frac{4}{\lambda_{\min}(\Sigma^{\pi})}\right)\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H\left(\Sigma^{\pi}\right)^{-1}\right)+2\sigma_{w}^{2}\left\|H\left(\Sigma^{\pi}\right)^{-1}\right\|\log\frac{1}{\delta}\right),$
where the final inequality follows from the fact that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$. This concludes the proof.
To guarantee concentration of the empirical covariance (25) to the true covariance, we state the following covariance concentration result.
@@
Let $\beta\in(0,\frac{\lambda_{\min}(\Sigma^{\pi})}{4})$. Suppose
$K\geq\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta))}{\beta^{2}}.$
@@
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta.$